OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = –2x + 8 (Points : 1) slope: 8, y-intercept: –2 slope: 2, y-intercept: –8 slope: –1, y-intercept: 4 slope: –2, y-intercept: 8

OpenStudy (anonymous):

@DarkBlueChocobo @cwrw238 @Kit_Kat_Nat_<3

OpenStudy (darkbluechocobo):

d

OpenStudy (anonymous):

that was fast lol

OpenStudy (darkbluechocobo):

slope is always x

OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = –x + 7 (Points : 1) slope: –7, y-intercept: 1 slope: –1, y-intercept: 7 slope: 1, y-intercept: –7 slope: 7, y-intercept: –1

OpenStudy (darkbluechocobo):

and its easy when its already in mx+b=y

OpenStudy (darkbluechocobo):

b

OpenStudy (anonymous):

David is selling floral arrangements. Each arrangement uses 1 vase and 12 roses. Each vase costs David $2.00. Let C be the total cost of the arrangement and r be the cost of 1 rose. Which equation should David use to find the total cost of each arrangement? (Points : 1) C = 12r + 2 C = 2r + 12 12C = r + 12 12 = C + 2r

OpenStudy (darkbluechocobo):

@Kit_Kat_Nat_<3 can you check these for me t-t im like 99% sure i just not sure

OpenStudy (anonymous):

maybe ask @cwrw238

OpenStudy (darkbluechocobo):

lelelel ive done this question before

OpenStudy (anonymous):

lol do u know the answer

OpenStudy (darkbluechocobo):

meh it just rings a bell :p aha

OpenStudy (darkbluechocobo):

so everything has to equal c

OpenStudy (darkbluechocobo):

which eliminates c and d

OpenStudy (anonymous):

so its either a or b?

OpenStudy (darkbluechocobo):

i think a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the slope of the line described by the equation 4x + 6y = 12?

OpenStudy (darkbluechocobo):

lelelel we just did this

OpenStudy (anonymous):

we did?

OpenStudy (darkbluechocobo):

yeh let me go double check the answer

OpenStudy (darkbluechocobo):

welp that one was -6y

OpenStudy (darkbluechocobo):

this is +

OpenStudy (darkbluechocobo):

4x + 6y = 12? 4x-12=-6y

OpenStudy (darkbluechocobo):

@wolf1728 would you divide this by 3? or 2?

OpenStudy (darkbluechocobo):

meh nvm i think divide whole thing by 6

OpenStudy (darkbluechocobo):

soh

OpenStudy (darkbluechocobo):

4/6x-2=y

OpenStudy (anonymous):

wait I'm confused what do u do after 4/6x -2 =y

OpenStudy (darkbluechocobo):

4/6x is your slope

OpenStudy (darkbluechocobo):

basicaly rise is 4 run is 6 if i am correct

OpenStudy (anonymous):

do u simplify is to 2/3? cause theres no 4/6 as a choice

OpenStudy (wolf1728):

y = –2x + 8 Basically, this is quite easy. Y intercept occurs when x = 0 y intercept = 8 The slope is the coefficient of x slope = -2

OpenStudy (anonymous):

there is no option for -2

OpenStudy (darkbluechocobo):

question wolf erm... How did you get +8 ?

OpenStudy (darkbluechocobo):

and still get rid of 6?

OpenStudy (darkbluechocobo):

it looks like you subtracted by 4?

OpenStudy (anonymous):

blue would it be 2/3?

OpenStudy (darkbluechocobo):

nah i screwed up somewhere let me investigate

OpenStudy (anonymous):

ok @MathematicsNerd

OpenStudy (darkbluechocobo):

4x + 6y = 12 4x-12=-6y 2/3x-2=-y -2/3x +2=y

OpenStudy (wolf1728):

y = -2x +8 y intercept occurs when x =0 y intercept = 0x +8 y intercept = 8

OpenStudy (darkbluechocobo):

@wolf1728 im still confused how you simplified the -6

OpenStudy (darkbluechocobo):

-6y*

OpenStudy (wolf1728):

Which equation are you working on?

OpenStudy (anonymous):

the last one i posted

OpenStudy (wolf1728):

y = -x + 7 Is this the one?

OpenStudy (darkbluechocobo):

finding the slope of this 4x + 6y = 12

OpenStudy (wolf1728):

4x + 6y =12 6y = -4x + 12 Dividing by 6 y = -(2/3)x +2 y intercept = -(2/3)0 + 2 y intercept = +2

OpenStudy (wolf1728):

Slope is the coefficient of x when y has been reduced to y with no coefficients. y = -(2/3)x +2 slope = -(2/3)

OpenStudy (darkbluechocobo):

Woop i was right then @wolf1728 thanks for confirming

OpenStudy (darkbluechocobo):

shaylee do you have -2/3? as answer to choose

OpenStudy (wolf1728):

Thanks DarkBlue

OpenStudy (darkbluechocobo):

@ShayleeCutie

OpenStudy (anonymous):

yes i got i thanks